OpenStudy (anonymous):

A package of spaghetti weighs 266 grams. How many ounces does the package weigh? 1 oz = 28 g

jimthompson5910 (jim_thompson5910):

1/28 = x/266 use the method described in the last problem to solve for x

jimthompson5910 (jim_thompson5910):

or you can say 1 oz = 28 g 1/28 oz = 1 g Keep going until you get 266 g = ??? oz